OpenStudy (rayford):

What are the x– and y–intercepts of y = 2x^2 – 10x + 8?

OpenStudy (lgbasallote):

x-intercept let y = 0 0 = 2x^2 - 10x + 8 now solve for x 2(x^2 - 5x + 4) = 0 x^2 - 5x + 4 = 0 (x-4)(x-1) = 0 x = 4; x = 1 <--x-intercepts y-intercept let x = 0 y = 2(0) - 10(0) + 8 y = 8 <--y-intrcept did you get that?

OpenStudy (rayford):

im sorry for the long pause..but yes i do :)

OpenStudy (lgbasallote):

wonderful